OpenStudy (anonymous):

Could someone please explain to me how to do this problem? Circle P has a radius of 8 units with center P at (2, -1). Which equation defines circle P? (x + 2)2 + (y – 1)2 = 64 (x – 2)2 + (y + 1)2 = 64 (x – 2)2 + (y + 1)2 = 8 (x + 2)2 + (y – 1)2 = 8

OpenStudy (anonymous):

equation 2

OpenStudy (phi):

the way you do this problem is memorize or look up the standard equation of a circle if you are given a center x0, y0 and a radius r, the equation is (x - x0)^2 + (y - y0)^2 = r^2

OpenStudy (anonymous):

so I just solve it out like algebra problems?

OpenStudy (phi):

to use the equation, replace the letters with the numbers they gave you (x - 2)^2 + (y - -1)^2 = 8*8 and simplify

OpenStudy (anonymous):

Oh okay Ill get back to you with my answer!

OpenStudy (phi):

notice you replace x0 with 2, y0 with -1, and r with 8

OpenStudy (phi):

the only tricky part is (y- -1)^2 becomes (y+1)^2

OpenStudy (phi):

and you have to know that r^2 means r*r so 8^2 is 8*8 or 64

OpenStudy (anonymous):

The general format for circle equation is: \[(x+x_{0})^{2}+(y+y_{0})^{2}=r ^{2}\] where r is the radius, 'X0' is the 'X' coordinate and 'Yo' is the 'Y' coordinate.

OpenStudy (phi):

Ghost, you have a typo: that should be x - x0 and y - y0 also, (x0,y0) are the coordinates of the center of the circle

OpenStudy (anonymous):

So I just realised I spent 10 minutes trying to solve it when I didnt need to haha. but thank you I get what I'm supposed to do. I cant medal you both so ghost medal phi and Ill medal you.

OpenStudy (anonymous):

\[(x-x_{0})^{2}+(y-y_{0})^{2}=r ^{2}\] knew I missed something, now better !